### v1

# Generalized Method of Moments

Abstract
We consider a generalized method of moments framework in which a part of the data vector is missing for some units in a completely unrestricted, potentially endogenous way. In this setup, the parameters of interest are usually only partially identified. We characterize the identified set for such parameters using the support function of the convex set of moment predictions consistent with the data. This identified set is sharp, valid for both continuous and discrete data, and straightforward to estimate. We also propose a statistic for testing hypotheses and constructing confidence regions for the true parameter, show that standard nonparametric bootstrap may not be valid, and suggest a fix using the bootstrap for directionally differentiable functionals of fang2019inference. A set of Monte Carlo simulations demonstrates that both our estimator and the confidence region perform well when samples are moderately large and the data have bounded supports.
Keywords: endogenous attrition, partial identification, panel data, support function
JEL codes: C14, C23
1 Introduction
Missing data are ubiquitous in empirical research, particularly in repeated-measurement settings. rubin1976inference classified mechanisms of missingness into three categories: missing completely at random (MCAR), missing at random (MAR), and missing not at random (MNAR). In this paper, we consider the last case, where missingness may be correlated with the data in an arbitrary manner. More specifically, we consider a generalized method of moments framework in which a portion of the data vector is missing for some units in a completely unrestricted, potentially endogenous way. A representative example is a two-period panel where, in the second period, some units drop out of the sample in a way that is correlated with both outcomes and covariates in both periods.
When no restriction on missingness is imposed, the parameters of interest are usually only partially identified. We characterize the identified set for the parameter of interest via the support function of the set of moments consistent with the observed data. We show that the identified set is sharp, is valid for both continuous and discrete data, and is straightforward to estimate.
We also show how to perform hypothesis testing and construct confidence regions using the estimate of the minimum of the support function as a test statistic. We derive the limit distribution of this test statistic and provide a procedure for computing critical values, following fang2019inference. We also show that the test controls size locally uniformly using an abstract result in fang2019inference.
Partial identification with missing data has been widely studied since Manski’s seminal work, Manski1989anatomy (see, e.g., Manski2005partial and Molinari2020microeconometrics for a survey). Our paper contributes to this extensive literature, with its key feature being that the identified set of moment predictions is convex, and the support function of this set can be estimated using a simple sample analog.
Our model is neither a moment inequality model with a convex identified set (e.g., KaidoSantos2014), nor the intersection bounds model of ChernozhukovLeeRosen2013intersection. It is similar to the model with convex moment predictions as in BeresteanuMolchanovMolinari2011, the main difference being that our characterization of the sharp identified set does not rely on a representation via random closed sets.
As mentioned above, the primary application of our model is a panel regression with attrition that may be endogenous. The analysis of panel data with attrition has a long history, and numerous techniques for handling attrition under various assumptions have been proposed. These include modelling the attrition process parametrically (e.g., HausmanWise1979), using auxiliary information in the form of refreshment samples (e.g., hirano2001combining), imposing semiparametric assumptions on the attrition process (e.g., bhattacharya2008inference), using inverse probability weighting (e.g., Wooldridge2002), selection models with unobserved heterogeneity (e.g., SemykinaWooldridge2010), and multiple imputation and Bayesian methods (e.g., DengEtAlSurvey).
The rest of the paper is organized as follows. Section 2 introduces the general framework and characterizes the sharp identified set. Section 3 suggests an estimator of the identified set. Section 4 develops a bootstrap procedure for hypothesis testing. Section 6 explores the performance of our methodology in a set of Monte Carlo simulations. Section 7 concludes. All proofs are given in the Appendix.
2 Setup and the identified set
We consider the generalized method of moments framework when some of the observed variables may be missing for some observations. Specifically, let $Z_{1i}$ be a data vector that we observe for each unit $i=1,\dots,n$ of the sample and let $Z_{2i}$ be a data vector that may be missing for some units. Let $S_{i}\in\{0,1\}$ be the sample selection indicator, which is equal to 1 when $Z_{2i}$ is observed and 0 otherwise. Dropping the subscript $i$, the observed vector is then $W=(S,Z_{1}^{\prime},SZ_{2}^{\prime})^{\prime}$. A key aspect of our framework is that we allow arbitrary dependence of $S$ on both observed and unobserved features. In particular, $S$ can be arbitrarily correlated with $Z_{2}$, representing endogenous sample selection. We do not allow for overlap between elements of $Z_{1}$ and $Z_{2}$. More generally, no features of the distribution of $Z_{2}$ are assumed to be identified from the random sample of $Z_{1}$. We are interested in a finite-dimensional parameter $\theta\in\Theta\subset\mathbb{R}^{d_{\theta}}$ that is defined by general moment conditions
$\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)]=0,$ (1)
where $\phi$ is a $d_{\phi}$-dimensional moment function and $\pi$ is the joint distribution of $Z_{1},Z_{2}$. Of course, $\pi$ is not point-identified due to the presence of missing data (see below), and hence neither is $\theta$.
One special case of this framework is two-period panel data with unrestricted attrition in the second period. To see this, let $Z_{t}=(X_{t}^{\prime},Y_{t})\in\mathbb{R}^{d}$ be the stacked vector of covariates and outcomes in period $t=1,2.$ In period 1, a random sample from $Z_{1}$ is observed. In period 2, there is attrition, and so $Z_{2}$ is only observed when $S=1$, where $S\in\{0,1\}$ is the indicator of staying in the sample. When the attrition is fully unrestricted and the parameter of interest is defined by a set of moment conditions (1) (such as, for instance, the slope coefficient in the fixed effects linear regression), this model becomes a special case of the GMM with missing data described above. For expositional clarity, we use the terminology related to this panel model throughout the rest of the paper.
We impose the following assumptions.
Assumption 1.
  1. (i)
    The parameter space $\Theta$ is compact.
  2. (ii)
    The distribution $\pi$ of $(Z_{1},Z_{2})$ belongs to the set $\mathcal{P}(\mathcal{Z}_{1}\times\mathcal{Z}_{2})$ of absolutely continuous distributions with respect to a known probability measure $\mu$ on $\mathcal{Z}_{1}\times\mathcal{Z}_{2}\subset\mathbb{R}^{2d}$, where $\mathcal{Z}_{1},\mathcal{Z}_{2}\subset\mathbb{R}^{d}$ are known compact sets.
  3. (iii)
    For each $\theta\in\Theta$, the map $(z_{1},z_{2})\mapsto\phi(z_{1},z_{2},\theta)$ is continuous.
  4. (iv)
    For each $\pi\in\mathcal{P}(\mathcal{Z}_{1}\times\mathcal{Z}_{2})$, the map $\theta\mapsto\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)]$ is continuous.
Assumptions 1(i), 1(iv) and 1(iii) are standard and impose compactness of the parameter space and the continuity of moments. Assumption 1(ii) is the requirement that the researcher knows which variables in the data are continuous and which are discrete. This knowledge determines the dominating measure $\mu$. Moreover, the researcher is assumed to be able to come up with a compact set $\mathcal{Z}_{1}\times\mathcal{Z}_{2}$ containing the support of the data. We impose no restrictions on the data-generating process beyond these basic requirements.
To characterize the identified set for $\theta$, let $p=P(S=1)\in(0,1)$ be the unconditional selection probability and notice that $\pi$ can be decomposed into the point-identified distribution of $(Z_{1},Z_{2})|S=1$, denoted by $\pi^{1}$, and the partially identified distribution of $(Z_{1},Z_{2})|S=0$, denoted by $\pi^{0}$, viz.,
$\pi=p\pi^{1}+(1-p)\pi^{0}.$
Denote by $\pi_{1}^{0}$ the point-identified probability distribution of $Z_{1}|S=0$. The sharp identified set for $\pi^{0}$ is the convex set of the distributions satisfying Assumption 1(ii) with first marginal $\pi_{1}^{0}$, i.e.,
$\Pi^{0}=\left\{\pi^{0}\in\mathcal{P}(\mathcal{Z}_{1}\times\mathcal{Z}_{2}):\,\,\pi^{0}(A\times\mathcal{Z}_{2})=\pi_{1}^{0}(A)\text{ for all Borel sets }A\subset\mathcal{Z}_{1}\right\}.$
Therefore, the sharp identified set for $\pi$ is
$\Pi=\left\{p\pi^{1}+(1-p)\pi^{0},\,\,\pi^{0}\in\Pi^{0}\right\}.$
Write
$\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)]=p\operatorname{\mathbb{E}}_{\pi^{1}}[\phi(Z_{1},Z_{2},\theta)]+(1-p)\operatorname{\mathbb{E}}_{\pi^{0}}[\phi(Z_{1},Z_{2},\theta)]=:\nu_{\pi^{0}}(\theta).$ (2)
Then the sharp identified set for $\theta$ is
$\Theta_{I}=\left\{\theta\in\Theta:\,\,\exists\pi^{0}\in\Pi^{0}\text{ such that }\nu_{\pi^{0}}(\theta)=0\right\}.$
The next proposition establishes a convenient characterization of $\Theta_{I}$ in terms of the support function of the compact, convex set [^2] See Step 1 in Appendix A for the proof of convexity and compactness of $N(\theta)$. of moment predictions
$N(\theta):=\left\{\nu_{\pi^{0}}(\theta):\,\,\pi^{0}\in\Pi^{0}\right\}.$
Proposition 1 (identified set via support function).
Suppose Assumption 1 holds and let
$\psi_{N(\theta)}(u)=\max_{\nu\in N(\theta)}u^{\prime}\nu$
be the support function of $N(\theta)$. Define the criterion function
$Q(\theta)=\min_{u\in\mathbb{B}}\psi_{N(\theta)}(u),$
where $\mathbb{B}$ is the unit ball in $\mathbb{R}^{d_{\phi}}$. Then $\Theta_{I}$ is compact and given by
$\Theta_{I}=\left\{\theta\in\Theta:\,\,Q(\theta)=0\right\}.$ (3)
Proof.
See Appendix A. ∎
Remark 2.1.
In the above characterization, the Euclidean unit ball  $\mathbb{B}$ can be replaced by any compact set containing the origin in its interior, such as the unit ball in the $\ell^{1}$ norm or the uniform norm. This may turn out to be convenient computationally.
Remark 2.2.
A similar strategy was recently employed by franguridi2025inference for characterizing the identified set in a partially identified GMM framework based on optimal transport.
The support function $\psi_{N(\theta)}$ can be written as
$\psi_{N(\theta)}(u)$ $=\max_{\pi\in\Pi}u^{\prime}\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)]$
$=p\operatorname{\mathbb{E}}[u^{\prime}\phi(Z_{1},Z_{2},\theta)|S=1]+(1-p)\max_{\pi^{0}\in\Pi^{0}}\operatorname{\mathbb{E}}_{\pi^{0}}[u^{\prime}\phi(Z_{1},Z_{2},\theta]].$
The last term involves optimization over an infinite-dimensional set of distributions. Fortunately, we can convert it into a finite-dimensional program for every value of $Z_{1}$, as the following proposition shows.
Proposition 2 (reduction formula).
Suppose Assumption 1 holds. Then
$\max_{\pi^{0}\in\Pi^{0}}\operatorname{\mathbb{E}}_{\pi^{0}}[u^{\prime}\phi(Z_{1},Z_{2},\theta)]=\operatorname{\mathbb{E}}\left[\max_{z_{2}\in\mathcal{Z}_{2}}u^{\prime}\phi(Z_{1},z_{2},\theta)|S=0\right].$
Proof.
See Appendix B. ∎
Figure 1: Set identification via support functions
Example 1.
Suppose there is only one moment condition, i.e., $\phi$ is scalar-valued. In this case, the set of moment predictions is the closed interval
$N(\theta)=\left[\min_{\pi\in\Pi}\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)],\,\,\max_{\pi\in\Pi}\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)]\right],$
The support function of $N(\theta)$ is a piecewise linear function given by
$\psi_{N(\theta)}(u)=\max_{\pi\in\Pi}u\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)]=\begin{cases}u\cdot\min_{\pi\in\Pi}\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)],&\text{if }u<0,\\ u\cdot\max_{\pi\in\Pi}\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)],&\text{if }u\geq 0,\end{cases}$
and hence the criterion function is
$Q(\theta)=\min_{u\in\mathbb{B}}\psi_{N(\theta)}(u)=\min\left\{0,\,\,\max_{\pi\in\Pi}\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)],\,\,-\min_{\pi\in\Pi}\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)]\right\}.$
The identification condition $Q(\theta)=0$ can be easily seen to be equivalent to
$\min_{\pi\in\Pi}\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)]\leq 0\leq\max_{\pi\in\Pi}\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)],$
which, of course, is just the condition that zero belongs to the set of moment predictions. See Figure 1 for an illustration of how the argmax set of $\psi_{N(\theta)}$ determines the position of $\theta$ relative to the identified set $\Theta_{I}$. Notice also that using the reduction formula in Proposition 2, we can write
$\min_{\pi\in\Pi}\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)]=\operatorname{\mathbb{E}}[f_{\min}(W,\theta)],$
$\max_{\pi\in\Pi}\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},\theta)]=\operatorname{\mathbb{E}}[f_{\max}(W,\theta)],$
where $W=(S,Z_{1}^{\prime},SZ_{2}^{\prime})^{\prime}$ is the data vector and
$f_{\min}(W,\theta)=S\phi(Z_{1},SZ_{2},\theta)+(1-S)\min_{z_{2}\in\mathcal{Z}_{2}}\phi(Z_{1},z_{2},\theta),$
$f_{\max}(W,\theta)=S\phi(Z_{1},SZ_{2},\theta)+(1-S)\max_{z_{2}\in\mathcal{Z}_{2}}\phi(Z_{1},z_{2},\theta).$
3 Estimation
Combining Propositions 1 and 2 suggests the following estimator of the identified set:
$\hat{\Theta}_{I}=\left\{\theta\in\Theta:\,\,\hat{Q}(\theta)\geq-\eta_{n}\right\},$
where $\eta_{n}>0$ is a tuning parameter and the sample criterion function
$\hat{Q}(\theta)=\min_{u\in\mathbb{B}}\hat{\psi}_{N(\theta)}(u)=\frac{1}{n}\sum_{i=1}^{n}\left(s_{i}u^{\prime}\phi(z_{i1},z_{i2},\theta)+(1-s_{i})\max_{z_{2}\in\mathcal{Z}_{2}}u^{\prime}\phi(z_{i1},z_{2},\theta)\right).$
Define the distance from a point $x\in\mathbb{R}^{d_{\phi}}$ to a set $A\subset\mathbb{R}^{d_{\phi}}$ by
$d(x,A)=\inf_{a\in A}\|a-x\|.$
For any sets $A,B\subset\mathbb{R}^{d_{\theta}}$, define the Hausdorff distance
$d_{H}(A,B)=\max\left\{\sup_{a\in A}d(a,B),\,\,\sup_{b\in B}d(b,A)\right\}.$
We impose the following assumptions.
Assumption 2.
There exists a weakly increasing function $m:\mathbb{R}_{+}\to\mathbb{R}_{+}$ such that $m(0)=0$, $m(\delta)>0$ for $\delta>0$ and $\min_{u\in\mathbb{B}}\psi_{N(\theta)}(u)\leq-m(d(\theta,\Theta_{I}))$.
Assumption 3.
$\eta_{n}\downarrow 0$ and $n^{-1/2}=o(\eta_{n})$.
Assumption 2 imposes separation of the identified set $\Theta_{I}$ in terms of the criterion function $Q(\theta)=\min_{u\in\mathbb{B}}\psi_{N(\theta)}(u)$. A sufficient condition can be formulated in terms of the gradient of $Q(\theta)$ away from the identified set. Assumption 3 states that $\eta_{n}$ has to converge to zero slower than $O(n^{-1/2})$.
The following proposition establishes the consistency of our estimator in the Hausdorff distance.
Theorem 1 (Hausdorff consistency).
Under Assumptions 1, 2 and 3, $d_{H}(\Theta_{I},\hat{\Theta}_{I})=o_{p}(1)$.
Proof.
See Appendix C. ∎
4 Inference
We now wish to develop a test of the hypothesis $H_{0}:\,\theta=\theta_{0}$. We suggest using the statistic
$\hat{T}(\theta_{0})=\sqrt{n}\min_{u\in\mathbb{B}}\hat{\psi}_{N(\theta_{0})}(u),$
which is the estimator of the scaled negative distance to the identified set of moment predictions $N(\theta_{0})$. Note that more negative values of $\hat{T}(\theta_{0})$ indicate stronger evidence against the null. To establish the asymptotic distribution of this statistic, we first prove that the function $\hat{\psi}_{N(\theta)}(u)$ converges to a Gaussian process uniformly over $u\in\mathbb{B}$ and $\theta\in\Theta$.
Assumption 4.
$\operatorname{\mathbb{E}}\left[\sup_{\theta\in\Theta}\sup_{z_{2}\in\mathcal{Z}_{2}}\|\phi(Z_{1},z_{2},\theta)\|^{2}\right]<\infty$.
Assumption 5.
There exists a random variable $L(Z_{1})$ such that $\operatorname{\mathbb{E}}L(Z_{1})^{2}<\infty$ and
$\sup_{z_{2}\in\mathcal{Z}_{2}}\|\phi(Z_{1},z_{2},\theta_{1})-\phi(Z_{1},z_{2},\theta_{2})\|\leq L(Z_{1})\|\theta_{1}-\theta_{2}\|$
for all $\theta_{1},\theta_{2}\in\Theta$, almost surely in $Z_{1}$.
Theorem 2 (uniform CLT for the support function).
Suppose Assumptions 1, 4 and 5 hold. Then there exists a Gaussian process $\mathbb{G}$ with trajectories in $C(\mathbb{B}\times\Theta)$ such that
$\sqrt{n}(\hat{\psi}_{N(\theta)}(u)-\psi_{N(\theta)}(u))\rightsquigarrow\mathbb{G}(u,\theta).$
Proof.
See Appendix D. ∎
Remark 4.1.
The results of KaidoSantos2014 suggest that the estimator $\hat{\psi}_{N(\theta)}(u)$ is asymptotically semiparametrically efficient for $\psi_{N(\theta)}(u)$.
An application of the functional delta method to Theorem 2 yields the following result for the asymptotic distribution of the test statistic $\hat{T}(\theta)$ over $\theta\in\Theta_{I}$.
Corollary 1 (asymptotic distribution of test statistic).
Under the assumptions of Theorem 2,
$\hat{T}(\theta)\rightsquigarrow\min_{u\in U(\theta)}\mathbb{G}(u,\theta)$
uniformly over $\theta\in\Theta_{I}$, where
$U(\theta)=\operatorname*{argmin}_{u\in\mathbb{B}}\psi_{N(\theta)}(u).$
Proof.
See Appendix E. ∎
Since the asymptotic distribution of $\hat{T}(\theta)$ is neither available in closed form nor is easy to simulate from, we need an alternative procedure for obtaining the critical values. The results of fang2019inference imply that, unless the Hadamard derivative $\min_{u\in U(\theta)}\mathbb{G}(u,\theta)$ is a linear functional of $\mathbb{G}$, the standard nonparametric bootstrap is invalid.
Example 1 (continued).
In the univariate case, the sample criterion function is
$\hat{Q}(\theta)=\min_{u\in\mathbb{B}}\hat{\psi}_{N(\theta)}(u)=\min\left\{0,\,\,\hat{\operatorname{\mathbb{E}}}[f_{\max}(W,\theta)],\,\,-\hat{\operatorname{\mathbb{E}}}[f_{\min}(W,\theta)]\right\}.$
Consider a data generating process and a parameter value $\theta$ satisfying $\operatorname{\mathbb{E}}[f_{\min}(W,\theta)]<0$, making the sample moment $\hat{\operatorname{\mathbb{E}}}[f_{\min}(W,\theta)]$ irrelevant asymptotically. We have
$Q(\theta)$ $=\min\left\{0,\,\,\operatorname{\mathbb{E}}[f_{\max}(W,\theta)]\right\},$
$\hat{Q}(\theta)$ $\overset{asy}{\sim}\min\left\{0,\,\,\hat{\operatorname{\mathbb{E}}}[f_{\max}(W,\theta)]\right\}.$
This is known to be an irregular problem when $\operatorname{\mathbb{E}}[f_{\max}(W,\theta)]=0$, which corresponds to $\theta$ being on the boundary of the identified set. In particular, the standard nonparametric bootstrap is invalid for the asymptotic distribution $\min(0,N(0,1))$ of $\sqrt{n}\hat{Q}(\theta)$, see, e.g., andrews2000inconsistency.
We show how to leverage the bootstrap for directionally differentiable functionals of fang2019inference. To this end, we propose an estimator of the derivative that satisfies their key Assumption 4.
For $\varepsilon_{n}>0$, define
$\hat{\chi}^{\prime}(h)=\min_{u\in\hat{U}(\varepsilon_{n})}h(u),$ (4)
where
$\hat{U}(\varepsilon_{n})=\left\{u\in\mathbb{B}:\,\,\hat{\psi}_{N(\theta_{0})}(u)\leq\min_{v\in\mathbb{B}}\hat{\psi}_{N(\theta_{0})}(v)+\varepsilon_{n}\right\}$
is the $\varepsilon_{n}$-enlarged” argmin set of $\hat{\psi}_{N(\theta_{0})}$. We impose the following assumptions.
Assumption 6 (sharp minima of $\psi_{N(\theta_{0})}$).
There exists $\kappa>0$ such that
$\psi_{N(\theta_{0})}(u)\geq\min_{v\in\mathbb{B}}\psi_{N(\theta_{0})}(v)+\kappa\cdot d(u,U(\theta_{0}))\text{ for all }u\in\mathbb{B}.$
Assumption 7 (bandwidth).
The sequence $\varepsilon_{n}$ satisfies $\varepsilon_{n}\downarrow 0$ and $\|\hat{\psi}_{N(\theta_{0})}-\psi_{N(\theta_{0})}\|_{\mathbb{B}}=o_{p}(\varepsilon_{n})$.
Assumption 6 posits that $\psi_{N(\theta_{0})}$ grows sufficiently fast away from its argmin set $U(\theta_{0})$. The latter does not have to be a singleton, as illustrated in Example 1 for which Assumption 6 holds. This assumption is equivalent to every subgradient $\nabla\psi_{N(\theta_{0})}(u)$ being bounded away from zero for $u\notin U(\theta_{0})$. It also suffices that it holds in a small neighborhood around $U(\theta_{0})$ rather than on the entire set $\mathbb{B}$. Assumption 7 restricts how fast $\varepsilon_{n}$ should converge to zero. Since $\|\hat{\psi}_{N(\theta_{0})}-\psi_{N(\theta_{0})}\|_{\mathbb{B}}=O_{p}(1/\sqrt{n})$ by Theorem 2, we can take $\varepsilon_{n}\sim\log n/\sqrt{n}$.
Given our estimator (4), the fang2019inference bootstrap algorithm for testing $H_{0}:\theta=\theta_{0}$ at the nominal size $\alpha\in(0,1)$ is as follows.
  1. 1.
    Compute $\hat{T}(\theta_{0})=\sqrt{n}\min_{u\in\mathbb{B}}\hat{\psi}_{N(\theta_{0})}(u)$.
  2. 2.
    Define $\hat{U}_{n}=\left\{u\in\mathbb{B}:\,\,\hat{\psi}_{N(\theta_{0})}(u)\leq\min_{v\in\mathbb{B}}\hat{\psi}_{N(\theta_{0})}(v)+\varepsilon_{n}\right\}$.
  3. 3.
    For each $b=1,\dots,B$:
    1. (a)
      Draw a random sample $W_{i}^{b}$, $i=1,\dots,n$, with replacement from the data $W_{i}=(S_{i},Z_{1i},S_{i}Z_{2i})$, $i=1,\dots,n$.
    2. (b)
      Compute $\hat{\psi}_{N(\theta_{0})}^{*b}(u)$ on the sample $W_{i}^{b}$, $i=1,\dots,n$.
    3. (c)
      Compute
      $\hat{T}^{*b}=\min_{u\in\hat{U}_{n}}\sqrt{n}\left(\hat{\psi}_{N(\theta_{0})}^{*b}(u)-\hat{\psi}_{N(\theta_{0})}(u)\right).$
  4. 4.
    Denote by $\hat{c}_{\alpha}^{*}$ the $\alpha$-quantile of $\hat{T}^{*1},\dots,\hat{T}^{*B}$.
  5. 5.
    Reject $H_{0}$ if $\hat{T}(\theta_{0})<\hat{c}_{\alpha}^{*}$.
The following theorem establishes that this bootstrap-based testing procedure controls size for a fixed data generating process.
Theorem 3 (test validity, pointwise).
Suppose Assumptions 6 and 7 hold and consider the testing procedure above with $B=\infty$. Then for every fixed data generating process, the asymptotic probability of rejecting the true null hypothesis is less than or equal to $\alpha$.
Proof.
See Appendix F. ∎
Although the previous theorem establishes that our test asymptotically controls size for a fixed data generating process, it is silent about uniform size control. In other words, it could be possible that even for a large sample size, one can find a particularly unfavorable DGP that would lead our test to exhibit significant size distortion. Such lack of uniform size control has been extensively studied in the context of moment inequality models, e.g., imbens2004confidence; romano2008inference; andrews2009validity; moon2009estimation; woutersen2006simple.
Fortunately, we can rely on the abstract results of fang2019inference to establish local size control. Denote by $P$ the joint distribution of the data vector $W=(S,Z_{1}^{\prime},SZ_{2}^{\prime})^{\prime}$ and let $\psi(P)$ be the corresponding population support function of $N(\theta_{0})$, where $\theta_{0}$ is any value in the identified set $\Theta_{I}=\Theta_{I}(P)$. We consider a sequence of DGPs $P_{t}$, $t\geq 0$, that is local to a given DGP $P_{0}$ in the following sense, cf. Assumption 5 in fang2019inference.
Assumption 8 (local sequence of DGPs).
The path $t\mapsto P_{t}$ is quadratic mean differentiable and satisfies for any $\lambda>0$ the following.
  1. (i)
    There exists $\psi^{\prime}(\lambda)$ such that
    $\|\sqrt{n}(\psi(P_{\lambda/\sqrt{n}})-\psi(P_{0}))-\psi^{\prime}(\lambda)\|_{\mathbb{B}}\to 0.$
  2. (ii)
    $\sqrt{n}(\hat{\psi}_{n}-\psi(P_{\lambda/\sqrt{n}}))\overset{\lambda}{\to}\mathbb{G}_{0}$, where $\overset{\lambda}{\to}$ denotes convergence in distribution under $\{W_{i}\}_{i=1}^{n}$ i.i.d., with each $W_{i}$ distributed as $P_{\lambda/\sqrt{n}}$, and $\hat{\psi}_{n}$ is our estimator of $\psi(P_{\lambda/\sqrt{n}})$ based on $\{W_{i}\}_{i=1}^{n}$.
  3. (iii)
    $\mathbb{G}_{0}$ is tight and supported on $C(\mathbb{B})$.
Assumption 8(i) imposes differentiability of the parameter $\psi(P)$ along the path $P_{\lambda/\sqrt{n}}$. It is expected to hold for any quadratic mean differentiable path in our model under mild regularity conditions. Assumption 8(ii) requires that the distribution of the estimator $\hat{\psi}_{n}$ is unaffected by local perturbations of the DGP, i.e. that $\hat{\psi}_{n}$ is regular. Finally, Assumption 8(iii) is expected to hold due to Theorem 2.
Theorem 4 (test validity, locally uniform).
Suppose Assumptions 6, 7 and 8 hold and consider the testing procedure above with $B=\infty$. Then this procedure controls size locally uniformly in the sense that
$\limsup_{n\to\infty}P_{\lambda/\sqrt{n}}\left(\hat{T}_{n}(\theta_{0})<\hat{c}_{\alpha}^{*}\right)\leq\alpha.$
Proof.
See Appendix H. ∎
5 Extension to more general missingness patterns
Our methodology can be extended to more general missingness patterns, as long as missingness is “monotone” (see below) and data contains no information on the distribution of the missing variables (i.e., missingness is “fully unrestricted”). We will illustrate this extension using the setup of panel data with more than two periods, where attrition happens monotonically, i.e., once the units drop out of the sample, they never return. For expositional simplicity, we consider the case of $T=3$ periods.
Denote the data in the three periods by $Z_{1},Z_{2},Z_{3}$ and denote the indicators of staying in the sample in periods 2 and 3 by $S_{2}$ and $S_{3}$, respectively. Since there is no return to the sample, $S_{2}=0$ implies $S_{3}=0$. Also, let
$p_{2}=P(S_{2}=1),\quad p_{3}=P(S_{3}=1),\quad p_{2|0}=p(S_{2}=1|S_{3}=0).$
The parameter of interest $\theta$ is defined by the moment conditions
$\operatorname{\mathbb{E}}_{\pi}[\phi(Z_{1},Z_{2},Z_{3},\theta)]=0,$
where the expectation is taken with respect to the latent, partially identified distribution $\pi$ of $(Z_{1},Z_{2},Z_{3})$.
Proposition 3.
The identified set is $\Theta_{I}=\{\theta\in\Theta:\,\,\min_{u\in\mathbb{B}}\psi_{\theta}(u)=0\}$, where
$\psi_{\theta}(u)$ $=p_{3}\operatorname{\mathbb{E}}[\phi(Z_{1},Z_{2},Z_{3},\theta)|S_{3}=1]+(1-p_{3})\left[p_{2|0}\operatorname{\mathbb{E}}[\max_{z_{3}}u^{\prime}\phi(Z_{1},Z_{2},z_{3},\theta)|S_{2}=1,S_{3}=0]\right$.
$+\left.(1-p_{2|0})\operatorname{\mathbb{E}}[\max_{z_{2},z_{3}}u^{\prime}\phi(Z_{1},z_{2},z_{3},\theta)|S_{2}=S_{3}=0]\right].$
Proof.
See Appendix G. ∎
6 Monte Carlo simulation
We consider a simple cross-sectional linear regression
$Y_{i}=\theta_{0}+\theta_{1}X_{i}+\varepsilon_{i},\quad i=1,\dots,n,$
where $Y_{i}$ is observed for the entire sample, whereas $X_{i}$ is missing for observations $i$ with $S_{i}=0$, where $S_{i}\in\{0,1\}$ is the sample selection indicator. The parameter of interest is $\theta=(\theta_{0},\theta_{1})$. We assume that the support of $X_{i}$ is known, in agreement with Assumption 1(ii). Dropping the subscript $i$, the moment conditions are
$\operatorname{\mathbb{E}}(Y-\theta_{0}-\theta_{1}X)$ $=0,$
$\operatorname{\mathbb{E}}[X(Y-\theta_{0}-\theta_{1}X)]$ $=0,$
and hence the support function is
$\psi_{\theta}(u_{1},u_{2})$ $=p\operatorname{\mathbb{E}}[u_{1}(Y-\theta_{0}-\theta_{1}X)+u_{2}X(Y-\theta_{0}-\theta_{1}X)\,|\,S=1]$
$+(1-p)\operatorname{\mathbb{E}}\left[\max_{x\in[0,1]}\left\{u_{1}(Y-\theta_{0}-\theta_{1}x)+u_{2}x(Y-\theta_{0}-\theta_{1}x)\right\}\right].$
Notice that the inner maximization is a constrained quadratic problem. We assume that $X_{i}$ has a uniform[0,1] distribution and $\varepsilon_{i}$ has a uniform[-1,1] distribution independently of $X_{i}$. Selection is completely random with probability $p=P(S_{i}=1)=0.9$. We set the sample size $n=1000$ and the tuning parameters $\eta_{n}=\varepsilon_{n}=0.1\log n/\sqrt{n}\approx 0.022$.
Figure 2: Estimates of the identified set and confidence regions averaged over 10001000 simulations.
The simulation results are illustrated in Figure 2. The true identified set (red contour) is approximated using a sample of $10,000$ observations. Our estimates of the identified set track the true identified set closely. The confidence regions are based on $1000$ bootstrap samples and are indicated by three black contour lines, one for each nominal coverage $1-\alpha\in\{0.10,0.90,0.95\}$. They are found to be quite narrow even at high nominal confidence levels.
7 Conclusion
In this paper, we study the partial identification of parameters defined by GMM moment conditions when components of the data vector are missing in a completely unrestricted manner. A leading motivation is panel data subject to endogenous attrition.
We characterize the sharp identified set through the support function of the convex set of moment predictions consistent with the observed data. For inference, we employ the minimum of the sample analog of the support function as a test statistic and develop a valid inference procedure, drawing on the bootstrap for directionally differentiable functionals of fang2019inference. We demonstrate that our estimator and confidence regions perform well in a set of Monte Carlo simulations.
[…]
Appendix A Proof of Proposition 1
Step 1: $N(\theta)$ is compact and convex.
Convexity is trivial. Since $\mathcal{Z}_{1}\times\mathcal{Z}_{2}$ is compact, Theorem 15.11 and Corollary 15.7 in aliprantis2006infinite imply that $\Pi^{0}$ is compact in the weak topology and the map $\pi^{0}\mapsto\nu_{\pi^{0}}(\theta)$ is continuous. Hence, $N(\theta)$ is compact as the image of a compact set under a continuous map.
Step 2: $\Theta_{I}=Q^{-1}(\{0\})$.
Since $N(\theta)$ is closed and convex, $0\in N(\theta)$ if and only if its support function is everywhere nonnegative: $\psi_{N(\theta)}(u)\geq 0$ for all $u\in\mathbb{R}^{\dim(\phi)}$. Indeed, if $0\in N(\theta)$, then $\psi_{N(\theta)}(u)=\max_{\nu\in N(\theta)}u^{\prime}\nu\geq 0$. Conversely, if $0\notin N(\theta)$, then by strong separation of a point from the closed convex set $N(\theta)$, there exists $u\neq 0$ and $\alpha>0$ such that $u^{\prime}\nu\leq\alpha$ for all $\nu\in N(\theta)$. This implies $\psi_{N(\theta)}(u)\leq-\alpha<0$.
Since $\psi_{N(\theta)}(0)=0$ for all $\theta$, nonnegativity of $\psi_{N(\theta)}$ is equivalent to the equality of
$\min_{u\in\mathbb{R}^{\dim(\phi)}}\psi_{N(\theta)}(u)$ (5)
to zero. Let us show that the latter is equivalent to the equality of
$\min_{u\in\mathbb{B}}\psi_{N(\theta)}(u)$ (6)
to zero. Indeed, if (5) is zero, then the minimum is achieved at $u=0$, and hence (6) is zero. Conversely, if (5) is nonzero, then there exists $u$ such that $\psi_{N(\theta)}(u)<0$. By the positive homogeneity of the support function, $\psi_{N(\theta)}(u/\|u\|)=\psi_{N(\theta)}(u)/\|u\|<0$, and hence the expression (6) is negative.
Step 3: $\Theta_{I}$ is compact.
Since $\Theta_{I}=Q^{-1}(\{0\})$, it suffices to establish the continuity of $Q(\theta)$. For this, notice that
$(u,\theta,\pi^{0})\mapsto u^{\prime}\nu_{\pi^{0}}(\theta)$
is a continuous function on the compact set $\mathbb{B}\times\Theta\times\Pi^{0}$, where $\Pi^{0}$ is equipped with the weak topology. Berge’s maximum theorem (see, e.g., Theorem 17.31 in aliprantis2006infinite) implies that the function
$(u,\theta)\mapsto\max_{\pi^{0}\in\Pi^{0}}u^{\prime}\nu_{\pi^{0}}(\theta)=\psi_{N(\theta)}(u)$
is continuous on the compact set $\mathbb{B}\times\Theta$. Applying Berge’s theorem again implies that $Q(\theta)$ is continuous, completing the proof.
Appendix B Proof of Proposition 2
Denote $c(z_{1},z_{2})=u^{\prime}\phi(z_{1},z_{2},\theta)$. Since $\pi^{0}\ll\mu$, each $\pi^{0}\in\Pi^{0}$ has a density $h=d\pi^{0}/d\mu$. Let
$\mu(dz_{1},dz_{2})=\mu_{1}(dz_{1})\,\mu(dz_{2}\mid z_{1})$
be the regular conditional decomposition of $\mu$. Then
$\pi^{0}(dz_{1},dz_{2})=h(z_{1},z_{2})\,\mu_{1}(dz_{1})\mu(dz_{2}\mid z_{1})$
and
$\int h(z_{1},z_{2})\,\mu(dz_{2}\mid z_{1})=\pi_{1}^{0}(z_{1})\qquad\text{for $\mu_{1}$-a.e.\ }z_{1}.$
For each $z_{1}$ with $\pi_{1}^{0}(z_{1})>0$, define
$g(z_{2}\mid z_{1}):=\frac{h(z_{1},z_{2})}{\pi_{1}^{0}(z_{1})}.$
Then $g(\cdot\mid z_{1})\geq 0$ and $\int g(z_{2}\mid z_{1})\,\mu(dz_{2}\mid z_{1})=1$, so $g(\cdot\mid z_{1})$ is a density with respect to $\mu(\cdot\mid z_{1})$. Therefore
$\operatorname{\mathbb{E}}_{\pi^{0}}[c(Z_{1},Z_{2})]=\int\pi_{1}^{0}(z_{1})\left(\int c(z_{1},z_{2})\,g(z_{2}\mid z_{1})\,\mu(dz_{2}\mid z_{1})\right)\mu_{1}(dz_{1}).$
Since any measurable family of densities $g(\cdot\mid z_{1})$ induces some $\pi^{0}\in\Pi^{0}$, maximizing over $\Pi^{0}$ reduces to choosing, for each $z_{1}$, a density $g(\cdot\mid z_{1})$ to maximize
$\int_{\mathcal{Z}_{2}}c(z_{1},z_{2})\,g(z_{2}\mid z_{1})\,\mu(dz_{2}\mid z_{1}).$
Fix $z_{1}$ and write $h_{z_{1}}(z_{2}):=c(z_{1},z_{2})$. Since $h_{z_{1}}$ is continuous on the compact set $\mathcal{Z}_{2}$, it attains its maximum at some $z_{2}^{\ast}(z_{1})$. For any density $g(\cdot\mid z_{1})$,
$\int h_{z_{1}}(z_{2})\,g(z_{2}\mid z_{1})\,\mu(dz_{2}\mid z_{1})\leq h_{z_{1}}(z_{2}^{\ast}(z_{1})).$
Moreover, for any $\varepsilon>0$ the continuity of $h_{z_{1}}$ implies the existence of an open neighborhood $U_{\varepsilon}$ of $z_{2}^{\ast}(z_{1})$ such that $h_{z_{1}}(z_{2})\geq h_{z_{2}^{\ast}(z_{1})}-\varepsilon$ for all $z_{2}\in U_{\varepsilon}$, and $\mu(U_{\varepsilon}\mid z_{1})>0$. Normalizing the indicator of $U_{\varepsilon}$ yields a density $g_{\varepsilon}(\cdot\mid z_{1})$ for which
$\int h_{z_{1}}(z_{2})\,g_{\varepsilon}(z_{2}\mid z_{1})\,\mu(dz_{2}\mid z_{1})\geq h_{z_{1}}(z_{2}^{\ast}(z_{1}))-\varepsilon.$
Letting $\varepsilon\downarrow 0$ shows that the supremum over densities equals the maximum value $h_{z_{1}}(z_{2}^{\ast}(z_{1}))$:
$\sup_{g(\cdot\mid z_{1})}\int c(z_{1},z_{2})\,g(z_{2}\mid z_{1})\,\mu(dz_{2}\mid z_{1})=\max_{z_{2}\in\mathcal{Z}_{2}}c(z_{1},z_{2}).$
Substituting this expression above yields
$\max_{\pi^{0}\in\Pi^{0}}\operatorname{\mathbb{E}}_{\pi^{0}}[c(Z_{1},Z_{2})]=\int\pi_{1}^{0}(z_{1})\,\max_{z_{2}\in\mathcal{Z}_{2}}c(z_{1},z_{2})\,\mu_{1}(dz_{1})=\operatorname{\mathbb{E}}\left[\max_{z_{2}\in\mathcal{Z}_{2}}c(Z_{1},z_{2})\,\middle|\,S=0\right],$
completing the proof.
Appendix C Proof of Theorem 1
We will follow the proof strategy in franguridi2025inference and show that $\Theta_{I}\subset\hat{\Theta}_{I}$ w.p.a. 1 and that for every $\delta>0$, $\hat{\Theta}_{I}\subset\Theta_{I}^{\delta}$ w.p.a. 1, where $\Theta_{I}^{\delta}$ is the $\delta$-enlargement of $\Theta_{I}$, i.e.
$\Theta_{I}^{\delta}=\left\{\theta\in\Theta:\,\,d(\theta,\Theta_{I})\leq\delta\right\}.$
Compactness of $\Theta_{I}$ will then imply the required Hausdorff convergence.
First, let us show that $\Theta_{I}\subset\hat{\Theta}_{I}$ w.p.a. 1. We have
$\sup_{\theta\in\Theta_{I}}\hat{Q}_{n}(\theta)\leq\sup_{\theta\in\Theta_{I}}Q(\theta)+\|\hat{Q}_{n}-Q\|_{\infty}=\|\hat{Q}_{n}-Q\|_{\infty}.$
By Theorem 2 and Assumption 3, the right-hand side is smaller than $\eta_{n}$ w.p.a. 1, and hence $\Theta_{I}\subset\hat{\Theta}_{I}$ w.p.a. 1.
Now let us show that for every $\delta>0$, $\hat{\Theta}_{I}\subset\Theta_{I}^{\delta}$ w.p.a. 1. By Assumption 2, we have $\inf_{\theta\notin\Theta_{I}^{\delta}}Q(\theta)\geq m(\delta)>0$. On the event $\|\hat{Q}_{n}-Q\|_{\infty}\leq r_{n}$, we have
$\inf_{\theta\notin\Theta_{I}^{\delta}}\hat{Q}_{n}(\theta)\geq\inf_{\theta\notin\Theta_{I}^{\delta}}Q(\theta)-\|\hat{Q}_{n}-Q\|_{\infty}\geq m(\delta)-r_{n}.$
Choose $n$ large enough so that $r_{n}\leq\eta_{n}<m(\delta)/2$. Then
$\inf_{\theta\notin\Theta_{I}^{\delta}}\hat{Q}_{n}(\theta)\geq m(\delta)/2>\eta_{n},$
and hence no $\theta$ outside of $\Theta_{I}^{\delta}$ belongs to $\hat{\Theta}_{I}$. Therefore, $\hat{\Theta}_{I}\subset\Theta_{I}^{\delta}$ w.p.a. 1.
Appendix D Proof of Theorem 2
Define the data vector $W=(S,Z_{1}^{\prime},SZ_{2}^{\prime})^{\prime}$ and let
$f_{u,\theta}(W)=Su^{\prime}\phi(Z_{1},SZ_{2},\theta)+(1-S)\max_{z_{2}\in\mathcal{Z}_{2}}u^{\prime}\phi(Z_{1},z_{2},\theta),$
so that $\hat{\psi}_{N(\theta)}(u)=\frac{1}{n}\sum_{i=1}^{n}f_{u,\theta}(W_{i})$ and $\psi_{N(\theta)}(u)=\operatorname{\mathbb{E}}f_{u,\theta}(W)$. For brevity, drop the arguments $Z_{1}$ and $SZ_{2}$ in $\phi(Z_{1},SZ_{2},\theta)$ and $\max_{z_{2}}\phi(Z_{1},z_{2},\theta)$. Write
$f_{u_{1},\theta_{1}}(W)-f_{u_{2},\theta_{2}}(W)$
$=S(u_{1}^{\prime}\phi(\theta_{1})-u_{2}^{\prime}\phi(\theta_{2}))+(1-S)\left(\max_{z_{2}}u_{1}^{\prime}\phi(z_{2},\theta_{1})-\max_{z_{2}}u_{2}^{\prime}\phi(z_{2},\theta_{2})\right)$
$=S\left((u_{1}-u_{2})^{\prime}\phi(\theta_{1})-u_{2}^{\prime}(\phi(\theta_{2})-\phi(\theta_{1}))\right)$
$+(1-S)\left(\max_{z_{2}}u_{1}^{\prime}\phi(z_{2},\theta_{1})-\max_{z_{2}}u_{2}^{\prime}\phi(z_{2},\theta_{1})+\max_{z_{2}}u_{2}^{\prime}\phi(z_{2},\theta_{1})-\max_{z_{2}}u_{2}^{\prime}\phi(z_{2},\theta_{2})\right).$
By Assumption 5, the term multiplied by $S$ can be bounded in absolute value by
$\|u_{1}-u_{2}\|\|\phi(Z_{1},SZ_{2},\theta_{1})\|+\|u_{2}\|L(Z_{1})\|\theta_{1}-\theta_{2}\|.$
For the first term multiplying (1S)(1-S), we have
$\left|\max_{z_{2}}u_{1}^{\prime}\phi(z_{2},\theta_{1})-\max_{z_{2}}u_{2}^{\prime}\phi(z_{2},\theta_{1})\right|\leq\max_{z_{2}}|(u_{1}-u_{2})^{\prime}\phi(z_{2},\theta_{1})|\leq\|u_{1}-u_{2}\|\max_{z_{2}}\|\phi(Z_{1},z_{2},\theta_{1})\|.$
Finally, for the second term multiplying (1S)(1-S), we have
$\left|\max_{z_{2}}u_{2}^{\prime}\phi(z_{2},\theta_{1})-\max_{z_{2}}u_{2}^{\prime}\phi(z_{2},\theta_{2})\right|\leq\max_{z_{2}}|u_{2}^{\prime}(\phi(z_{2},\theta_{1})-\phi(z_{2},\theta_{2}))|$
$\leq\|u_{2}\|\max_{z_{2}}\|\phi(z_{2},\theta_{1})-\phi(z_{2},\theta_{2})\|\leq L(Z_{1})\|\theta_{1}-\theta_{2}\|,$
where the last inequality follows from $\|u_{2}\|\leq 1$ and Assumption 5.
Combining the inequalities above with Assumption 4 implies that the function class $\mathscr{F}=\{f_{u,\theta}:\,u\in\mathbb{B},\,\theta\in\Theta\}$ is Lipschitz in the parameter with a square integrable constant. By Theorem 2.7.17 in vaart2023empirical, $\mathscr{F}$ is Donsker.
Appendix E Proof of Corollary 1
For any $\psi\in C(\mathbb{B})$, the functional $\psi\mapsto\min_{u\in\mathbb{B}}\psi(u)$ is Hadamard directionally differentiable with the derivative
$h\mapsto\min_{u\in U(\psi)}h(u),\quad h\in C(\mathbb{B}),$
where
$U(\psi)=\arg\min_{u\in\mathbb{B}}\psi(u),$
as established in Theorem 3.1 in shapiro1991asymptotic. Applying the functional delta method (e.g., Theorem 3.10.5 in vaart2023empirical) to Theorem 2 completes the proof.
Appendix F Proof of Theorem 3
We need the following two technical lemmas, which are completely abstract and independent of our framework.
Lemma 1.
Suppose Assumptions 6 and 7 hold. Let $\hat{\psi}$ be random functions on $\mathbb{B}$ converging uniformly in probability to $\psi$. Then
$d_{H}(\hat{U}(\varepsilon_{n}),U)=o_{p}(1).$
Proof.
The Hausdorff convergence is implied by
$\sup_{u\in\hat{U}(\varepsilon_{n})}d_{H}(u,U)=o_{p}(1)$ (7)
and
$\sup_{u\in U}d_{H}(u,\hat{U}(\varepsilon_{n}))=o_{p}(1).$ (8)
First, let us establish (7). Let $\delta_{n}=\|\hat{\psi}-\psi\|_{\mathbb{B}}$. For $v\in\hat{U}(\varepsilon_{n})$, we have
$\psi(v)\leq\hat{\psi}(v)+\delta_{n}\leq\min_{u\in\hat{U}(\varepsilon_{n})}\hat{\psi}(u)+\varepsilon_{n}+\delta_{n}\leq\min_{u\in U}\psi(u)+2\delta_{n}+\varepsilon_{n},$
where the second inequality holds by the definition of $\hat{U}(\varepsilon_{n})$. Combining with Assumption 6 yields
$\kappa\cdot d_{H}(v,U)\leq\psi(v)-\min_{u\in U}\psi(u)\leq 2\delta_{n}+\varepsilon_{n}.$
Taking the supremum over $v\in\hat{U}(\varepsilon_{n})$ and noting that $2\delta_{n}+\varepsilon_{n}\to 0$ establishes (7).
Now let us establish (8). Indeed, for any $u\in U$,
$\hat{\psi}(u)\leq\psi(u)+\delta_{n}=\min_{u\in U}\psi(u)+\delta_{n}\leq\min_{u\in\hat{U}(\varepsilon_{n})}\hat{\psi}(u)+2\delta_{n}.$
Consider the event $E_{n}=\{\delta_{n}\leq\varepsilon_{n}/2\}$. Using the inequality above, on this event,
$\hat{\psi}(u)\leq\min_{u\in\hat{U}(\varepsilon_{n})}\hat{\psi}(u)+\varepsilon_{n},$
which implies $u\in\hat{U}(\varepsilon_{n})$, and hence $U\subset\hat{U}(\varepsilon_{n})$. By Assumption 7, $P(E_{n})\to 1$, and so $U\subset\hat{U}(\varepsilon_{n})$ w.p.a. 1. Therefore, $\sup_{u\in U}d_{H}(u,\hat{U}(\varepsilon_{n}))=0$ w.p.a. 1, establishing (8). ∎
Lemma 2.
Suppose $d_{H}(\hat{U},U)=o_{p}(1)$, where $\hat{U},U$ are nonempty compact subsets of $\mathbb{B}$. Then for any continuous function $h$ on $\mathbb{B}$,
$\min_{u\in\hat{U}}h(u)-\min_{u\in U}h(u)=o_{p}(1).$
Proof.
The modulus of continuity of $h$ is
$\omega_{h}(\varepsilon)=\max_{u_{1},u_{2}\in\mathbb{B}:\,\|u_{1}-u_{2}\|\leq\varepsilon}|h(u_{1})-h(u_{2})|.$
Since $h$ is continuous and $\mathbb{B}$ is compact, $\omega_{h}(\varepsilon)\downarrow 0$ as $\varepsilon\downarrow 0$.
Pick $u^{*}\in\arg\min_{u\in U}h(u)$. By definition of the Hausdorff distance, there exists $\hat{u}\in\hat{U}$ such that $\|\hat{u}-u^{*}\|\leq d_{H}(\hat{U},U)$. Then
$\min_{u\in\hat{U}}h(u)\leq h(\hat{u})\leq h(u^{*})+\omega_{h}(d_{H}(\hat{U},U)),$
which implies
$\min_{u\in\hat{U}}h(u)-\min_{u\in U}h(u)\leq\omega_{h}(d_{H}(\hat{U},U)).$
Similarly, pick any $\hat{u}^{*}\in\arg\min_{u\in\hat{U}}h(u)$. There exists $u\in U$ such that $\|\hat{u}^{*}-u\|\leq d_{H}(\hat{U},U)$. Then
$\min_{u\in U}h(u)\leq h(\hat{u}^{*})+\omega_{h}(d_{H}(\hat{U},U)),$
which implies
$\min_{u\in U}h(u)-\min_{u\in\hat{U}}h(u)\leq\omega_{h}(d_{H}(\hat{U},U)).$
But since $\omega_{h}(\varepsilon)$ is continuous at $\varepsilon=0$ and $d_{H}(\hat{U},U)=o_{p}(1)$, by the continuous mapping theorem, $\omega_{h}(d_{H}(\hat{U},U))=o_{p}(1)$, completing the proof. ∎
Let us show that the estimator $\hat{\chi}^{\prime}$ is amenable to Remark 3.4 in fang2019inference. Indeed, let $\hat{\Delta}=\|h_{1}-h_{2}\|_{\hat{U}(\varepsilon_{n})}$ (uniform norm on $\hat{U}(\varepsilon_{n})$). Then $h_{2}(u)\geq h_{1}(u)-\hat{\Delta}$, and so
$\min_{u\in\hat{U}(\varepsilon_{n})}h_{2}(u)\geq\min_{u\in\hat{U}(\varepsilon_{n})}h_{1}(u)-\hat{\Delta}$.
Swapping $h_{1}$ and $h_{2}$ yields the reverse inequality. Combining the two inequalities yields
$\left|\hat{\chi}^{\prime}(h_{1})-\hat{\chi}^{\prime}(h_{2})\right|=\left|\min_{u\in\hat{U}(\varepsilon_{n})}h_{1}(u)-\min_{u\in\hat{U}(\varepsilon_{n})}h_{2}(u)\right|\leq\hat{\Delta}\leq\|h_{1}-h_{2}\|_{B},$
i.e. $\hat{\chi}^{\prime}$ is 1-Lipschitz. Therefore, according to Remark 3.4, it suffices to establish pointwise consistency of $\hat{\chi}^{\prime}(h)$, i.e., for any $h\in C(\mathbb{B})$,
$\left|\hat{\chi}^{\prime}(h)-\chi^{\prime}(h)\right|=o_{p}(1).$
Indeed, Theorem 2 implies that $\hat{\psi}$ converges uniformly in probability to $\psi$ on $\mathbb{B}$. By Lemma 1, the argmin sets $\hat{U}(\varepsilon_{n})$ converge to $U$ in the Hausdorff distance in probability. Using Lemma 2 completes the proof.
Appendix G Proof of Proposition 3
Dropping the argument $\theta$ for brevity,
$\operatorname{\mathbb{E}}[\phi(Z_{1},Z_{2},Z_{3})]$ $=p_{3}\operatorname{\mathbb{E}}[\phi(Z_{1},Z_{2},Z_{3})|S_{3}=1]+(1-p_{3})\operatorname{\mathbb{E}}[\phi(Z_{1},Z_{2},Z_{3})|S_{3}=0].$
Notice that $\operatorname{\mathbb{E}}[\phi(Z_{1},Z_{2},Z_{3})|S_{3}=1]$ is point identified since $S_{3}=1$ implies $S_{2}=1$. Then
$\operatorname{\mathbb{E}}[\phi(Z_{1},Z_{2},Z_{3})|S_{3}=0]$ $=p_{2|0}\operatorname{\mathbb{E}}[\phi(Z_{1},Z_{2},Z_{3})|S_{2}=1,S_{3}=0]$
$+(1-p_{2|0})\operatorname{\mathbb{E}}[\phi(Z_{1},Z_{2},Z_{3})|S_{2}=S_{3}=0].$
Arguing as in the proof of Proposition 2 completes the proof.
Appendix H Proof of Theorem 4
We map our testing framework into Section 3.4.2 of fang2019inference and then apply their Corollary 3.2. To this end, multiply our statistic by 1-1 to obtain
$T(\theta)$ $=\chi(\psi_{\theta})=-\min_{u\in\mathbb{B}}\psi_{\theta}(u),$
$\hat{T}(\theta)$ $=\chi(\hat{\psi}_{\theta})=-\min_{u\in\mathbb{B}}\hat{\psi}_{\theta}(u),$
where
$\chi(h)=-\min_{u\in\mathbb{B}}h(u).$
This functional is Hadamard differentiable at $\psi_{\theta}$ with the derivative
$\chi_{\psi_{\theta}}^{\prime}(h)=-\min_{u\in U}h(u),$
where $U=\arg\min_{u\in\mathbb{B}}\psi_{\theta}(u)$. Notice that $T(\theta)=0$ if $\theta\in\Theta_{I}$ and $T(\theta)>0$ otherwise. Hence, we are testing the hypothesis $H_{0}:\chi(\psi_{\theta_{0}})\leq 0$ against $H_{1}:\chi(\psi_{\theta_{0}})>0$, as in equation (31) of fang2019inference.
Let us show now that the functional $\chi^{\prime}(h)$ is convex. Indeed, for all $h_{1},h_{2}$ and $u\in U$, we have
$th_{1}(u)+(1-t)h_{2}(u)\geq t\min_{u\in U}h_{1}(u)+(1-t)\min_{u\in U}h_{2}(u),$
and hence
$-\chi^{\prime}(th_{1}+(1-t)h_{2})$ $=\min_{u\in U}\left\{th_{1}(u)+(1-t)h_{2}(u)\right\}$
$\geq t\min_{u\in U}h_{1}(u)+(1-t)\min_{u\in U}h_{2}(u)=-t\chi^{\prime}(h_{1})-(1-t)\chi^{\prime}(h_{2}).$
Multiplying this inequality by 1-1 establishes the convexity of $\chi^{\prime}(h)$. Applying Corollary 3.2 in fang2019inference completes the proof.